OpenStudy (anonymous):

calculate the moments mx and my and the center of mass of a lamina with the given density and shape. P = 3 (density) y= (1 -x^2)^1/2 from x = 0 to 1 y = x - 1 from x = 0 to 1

OpenStudy (anonymous):



OpenStudy (anonymous):

what seem to be the problem, you know the formulas and everything right

OpenStudy (anonymous):

i need to find the centroid

OpenStudy (anonymous):

okay, but you know the formulas and everything right

OpenStudy (anonymous):

yeah so would i use f(x) = \[\sqrt{1 - x ^{2}}\] and g(x) = x - 1?

OpenStudy (anonymous):

yes then simply have to take into account that density

OpenStudy (anonymous):

for the formula y bar = 1/a\[\int\limits_{0}^{1}(f ^{2}x - g ^{2}x)?\]

OpenStudy (anonymous):

your squareing both function why? to get rid of the root

OpenStudy (anonymous):

well thats the formula my teacher told me for y bar

OpenStudy (anonymous):

but then i get my = 3\[\int\limits_{0}^{1}x/2(\sqrt{1 - x ^{2}} -x + 1)\] and idk how to do the integral of that